At s.t.p,a certain Mass of gas occupies a volume of 800cm3. find the temperature at which the glass occupies 1000cm3 and has a pressure of 720mmhg.
1
Expert's answer
2021-08-23T02:56:47-0400

At S.TP

Pressure = 760mmHg

Temperature = 273


"\frac{P_1V_1}{T_1}=
\frac{P_2V_2}{T_2}"


"T_2=" "\frac{P_2V_2T_1}{P_1V_1}"


"T_2=\frac{720×1000×273}{760×800}=323.29K"